Atsie (atsie):

Truly trying to get the hang of this! :) Check me if I'm right or wrong and explain yay or nay. ^.^

Atsie (atsie):

I'm thinking y = -4x + 2 Because -2 minus 2 is -4 which divided by 2 is -2. I'm all over the universe, but some explanation would be nice! :))

pooja195 (pooja195):

Your answer is correct nice job

Atsie (atsie):

It is!?? I'm totally shocked! Thank you Pooja :)

MARC (marc_d):

Step 1:use this slope formula to find the value of slope,m\[\huge{m=\frac{ y_2-y_1 }{ x_2-x_1 }}\]let say point\[(0,2)=(x_1,y_1)\]and point\[(1,-2)=(x_2,y_2)\]Then,substitute the value into the slope formula\[m=\frac{ -2-2 }{ 1-0 }\]\[m=\frac{ -4 }{ 1 }\]\[m=-4\] Okay,now we have the value of the slope which is -4 and the value of c(y-intercept) which is 2 Step 2:Use this formula to form the equation\[\huge{y=mx+c}\]where m=slope c=y-intercept okay..now we hv the value of m and c. substitute the value of the slope and the value of c(y-intercept)into the formula y=mx+c\[\huge{y=\color{red}{m}x+\color{red}{c}}\]